# supplement Analyzing Defect Complex Structures of PL6 in SiC: Insights from First-Principles Screening

### v1

@@ Abstract
The PL6 color center in 4H-SiC has attracted significant research attention owing to its remarkable spin and optical properties at ambient temperatures. Unlike many color centers that require cryogenic conditions to function, the PL6 center maintains optical activity even at room temperature or higher, benefiting from the material’s high thermal conductivity and chemical stability. To elucidate the origin of PL6’s distinctive characteristics and clarify its microscopic structure in the crystal, this study conducts a comprehensive investigation of potential defect complexes through systematic first-principles calculations. Our computational approach rigorously evaluates the electronic structures and spin-related properties of various defect configurations, enabling reasonable identification of the most probable atomic arrangement responsible for the observed characteristics of PL6. Our analysis identifies a limited set of candidate structures with -like symmetry consistent with known experimental signatures. Among these, two defect complexes emerge as the most likely PL6 candidates. The findings suggest that targeted experimental studies, such as electron spin resonance and optically detected magnetic resonance, could further clarify the nature of the PL6 center, while additional theoretical investigations on transition matrix elements and defect formation dynamics are warranted.
@@ I Introduction
Spin defects in solid-state systems are promising candidate systems for quantum information science [1]. Among these candidates, divacancies in 4H-SiC have garnered significant attention due to their long coherence time [2, 3], spin-selective optical transitions [4], and telecom-wavelength fluorescence [5]. Photoluminescence (PL) centers, specifically PL1-8 [6, 7], have been reported in 4H-SiC. In particular, PL5 and PL6 exhibit similar optical and spin characteristics to those of PL1-4 [8, 6] but exhibit greater superior stability compared to PL1-4 under optical excitation [9, 10]. Experiment [8] and first-principles calculations [11] have unambiguously identified PL1-4 as the hh, kk, hk, and kh divacancies, but PL5-8 centers have not been fully identified.
Due to their highly similar properties, PL5 and PL6 are generally considered to be perturbed versions of the PL1-4 divacancy defects, influenced by their surrounding environment. One proposed explanation for the distinct properties of PL5 and PL6 related them to kh and kk divacancies located near stacking faults [12], while, challenged by subsequent experiments [13]. Divacancies generated via post-irradiation annealing (following electron or ion bombardment) [14] exist within a multifaceted point defect landscape in SiC, coexisting with intrinsic vacancies, antisite defects, and potential residual impurities inherent in the as-grown SiC material. Given this complex defect landscape, another plausible hypothesis emerges: PL5 and PL6 are complexes formed from a divacancy paired with another intrinsic point defect. In support of this concept, a recent study combining experimental and theoretical results proposed a new atomic model for PL5. This model identifies PL5 as a kh divacancy complexed with a nearby carbon antisite () defect [15]. Adopting this composite defect perspective – that PL6, like PL5, also involves a divacancy and another intrinsic defect – our current work aims to identify the possible atomic structures of PL6. We employ first-principles calculations within a constructed selection scheme to screen potential defect complex structures for PL6. This screening is based on zero-field splitting (ZFS), estimation of optical transition energy based on the Kohn-Sham (KS) gap, and binding energy of compounded defects. Ultimately, we identify two defect complex structures as the most probable candidates for PL6.
@@ II.1 Calculation setup
II.2 Determining Candidate Compounded Defects
Experimental observations show that the rhombic ZFS parameter of PL6 is almost zero [6, 8], suggesting the principal axis of PL6 is parallel to the c-axis. Consistent with this, angle-resolved measurements also indicate a -like symmetry for PL6 [13]. Based on these experimental findings and our prior considerations, hh and kk are proposed as parent models for PL6. To search for the possible defect complex structure as a variant of hh or kk, a cubic-like supercell with basis vectors (where are the basis vectors of the primitive cell) was constructed. In the first stage, a hh or kk defect was generated, with its middle placed at the center of the supercell. A cylinder with a height of 8 Å and a base radius of 7 Å centered at the defect center was chosen as the total selection range for candidates. Next, the symmetry of hh and kk was used to exclude redundant atom sites in the cylinder structure. For the remaining sites, we only consider the neutral antisite defect, i.e., the selection is conducted on the neutral complex defect consisting of hh/kk and antisite). The final selection range contains 116 cases, 56 for hh and 60 for kk. Our selection range covers all the antisites that may have a significant impact on the divacancies properties, as the spin density distribution of the defect does not uniformly decay with distance from the hh/kk center (see Fig. LABEL:spin-density in the Supplemental Material [23]).
PL6 exhibits an ground state with overall similar but distinct properties compared to PL1-4. This implies that the electron numbers of PL6 must be even, and its electronic configuration should resemble that of divacancy defects. Crucially, introducing an additional vacancy would lead to a state with , thereby definitively excluding vacancy-related configurations from consideration. Regarding antisite defects, since and act as isovalent substituents, a neutral antisite defect combined with a divacancy would alter the local structure but maintain a similar electronic environment, thus preserving the fundamental spin and optical properties. Although the charge state also satisfies the even electron number requirement, its electronic configuration significantly deviates from the divacancy case. Consequently, this study focuses exclusively on the hh/kk-type defect complexes incorporating neutral antisite defects.
@@ II.3 Screening Methodology for the PL6 Center Identification
To systematically identify the microscopic origin of the PL6 center in SiC, we established a multi-criteria screening framework that synergistically combines spin, optical, and thermodynamic properties. The selection criteria were strategically chosen to balance computational feasibility with experimental relevance, prioritizing three key aspects: (1) Consistency with experimentally measured ZFS parameters; (2) Compatibility with photoluminescence characteristics through KS level analysis; (3) Thermodynamic stability of defect complexes. This multi-property approach enables effective discrimination among thousands of possible defect configurations while maintaining manageable computational costs.
@@ II.3.1 ZFS Parameter Calibration and Error-Tolerant Thresholds
Note that experimentally measured values correspond to their absolute magnitudes. The ZFS parameters (, ) serve as critical fingerprints for spin defects. To address systematic errors in our DFT implementation (including PAW correction and spin decontamination errors), we performed benchmark calculations on the four well-characterized divacancies (PL1-4, i.e., hh, kk, hk, kh defects), as shown in Table 1. The calculated values are 10% systematically overestimated and the calculated values have fluctuations compared to experimental data. To ensure comprehensive coverage of potential PL6 candidates while accounting for these errors, we established conservative thresholds:
[1.365, 1.706] GHz (1.0–1.25 experimental = 1.365 GHz)
20 MHz (experimental = 0 MHz ± 20 MHz)
@@
| (GHz) | (MHz) | |||
| Divacancies | expt | calc | expt | calc |
@@ II.3.2 KS Level Difference as ZPL Predictor in Defect Screening
The photoluminescence characteristics provide critical complementary signatures for defect identification. For the symmetry divacancies, the ZPL originates from the spin-allowed transition [25], corresponding to two specific occupations of KS orbitals differed by promoting an electron from the orbital to the orbital in the minority spin channel. The calculation of ZPL of PL1-4 has been detailed and tested among supercell size, k-point mesh, and functional selection [11]. However, such calculations are computationally consumption (e.g., HSE06 functional, a -only calculation with a supercell of 2400 atoms, or a calculation with -mesh for a 96-atom supercell) [11, 26].
The observed redshift in ZPL of PL6 over the hh/kk divacancies [8, 9] suggests a perturbation of the splitting between KS orbitals from an antisite defect. We establish a computationally efficient proxy for ZPL shifts through KS level analysis: the energy difference between the highest occupied () and lowest unoccupied () minority spin states systematically correlates with experimental ZPL trends in benchmarked defects (PL1-4). This relationship arises because ZPL energies and KS splitting respond proportionally to the perturbation. By requiring candidate complexes to exhibit enhanced KS splitting compared to isolated hh/kk centers, we effectively screen for structures capable of reproducing PL6’s characteristic ZPL shift.
@@ II.3.3 Binding Energy for Stability of Complex Defects
where is the formation energy of point defects. Negative values indicate exothermic complex formation, serving as a necessary (though not sufficient) condition for stability. While full thermodynamical analysis would require considering formation pathways [28] and configurational entropy [29], our screening conservatively requires to filter physically plausible complexes.
As a summary, based on experimental measurements yielding GHz and MHz for the PL6 center, we define the selection criteria for its spin properties as follows: within GHz and within MHz. To account for the observed ZPL blue shift of PL6 relative to hh and kk, we require PL6’s KS gap to exceed those of its reference hh/kk. Additionally, a negative binding energy is imposed to ensure the stability of the defect structure. Despite the intentionally permissive criteria, the geometric and electronic constraints collectively yield only a few configurations, demonstrating remarkable latent selectivity in defect screening.
@@ III Results and Discussion
III.1 The Possible PL6 Structure
After applying the screening strategy to the 116 candidates, only 12 configurations remain, as listed in Table 2. To systematically characterize these surviving configurations, the unperturbed divacancy, and associated antisite defects are labeled by a grid-based positional encoding scheme. The numerical indices following an antisite symbol specify its position relative to the of the divacancy using grid coordinates along the -, -, and -axes, where negative displacements are denoted by overlined numbers. These coordinate axes are defined according to a gridded supercell with spacings of 1.540 Å, 0.889 Å, and 0.628 Å along each respective axis. Complete implementation details with visual examples are provided in the Supplemental Material [23].
@@ III.1 The Possible PL6 Structure
A notable feature of the results is that all the binding energies are relatively weak, but for , with a value of eV, while all the others are at or lower than eV. This indicates that for most candidates, the attraction between the divacancy and antisite defect is weak, leading to defect complexes with limited stability. Another feature is that most of the candidates possess an whose absolute value is in the range of 10 to 20 MHz. A more stringent screening criterion on E (absolute value within 10 MHz) thus yields the following candidates: , , and . Additionally, the remaining candidates show weak KS gap enhancement relative to the original kk divacancy defect, which is consistent with the small binding energy. Among these candidates, the most enhanced KS level splitting is and . Considering all these aspects, and are the most likely PL6 candidates.
@@
Apart from the difference in local environment between and , the dominant difference between them is the distance between the kk center and the antisite, which is 4.10 Å for and 6.00 Å for . The longer distance in leads to a weaker interaction between the kk and the antisite, which is consistent with its smaller binding energy. For these two defects, we further calculated the actual ZPL energy by using constrained DFT calculation and the 0/1- charge transition level by the PBEsol functional. While this set of computational parameters cannot reliably reproduce experimental ZPL values, it can still capture relative shifts between different configurations. This allows for a comparative analysis of trends despite the absolute discrepancy. A comparison of the results with kk is shown in Table 3. shows a larger ZPL shift of about 0.13 eV than kk, while the ZPL shift of is only about 0.03 eV. Moreover, the 0/-1 charge transition level of is closer to the CBM than kk, consistent with the experimental charge dynamics model [9]. Thus, though has a relatively small binding energy, it is the most possible PL6 candidate. Although the is closer to the kk center in , the properties of the divacancy mainly arise from the dangling bond of the C atom, which originates from the inside the divacancy. Although is farther from the kk center, it is closer to the dangling bond of the kk center, which is the reason why the impact of the antisite defect on the kk center is larger in than in .

### v2

# supplement_aux Towards Identifying the PL6 Center in SiC: From First-Principles Screening to Hyperfine Validation of Competing Defect Candidates

@@ Abstract
The PL6 color center in 4H-SiC, known for its excellent ambient-temperature spin and optical properties, has an unresolved microscopic origin. In this first-principles study, we systematically investigate potential structures to clarify its nature. We first rigorously examine the DV-antisite hypothesis (a divacancy paired with a carbon antisite, ), analyzing the energetic, electronic, and spin properties of various configurations. Two -symmetric complexes emerge as strong candidates within this framework. Subsequently, a critical comparison of hyperfine interaction signatures is performed between these candidates, the alternative OV model [specifically OV(hh) and OV(kk), an oxygen replacing C together with a Si vacancy], and experimental data. This analysis demonstrates that the OV(hh) structure more accurately reproduces PL6’s hyperfine features. Furthermore, re-evaluation of the proposed OV(hk) model for the related PL5 center reveals zero-field splitting parameter inconsistencies with experimental results, suggesting that PL5 and PL6 may have distinct origins. These findings provide crucial theoretical insights and motivate targeted experimental validation for these quantum defects.
@@ I Introduction
Spin defects in solid-state systems are promising candidate systems for quantum information science [1]. Among these candidates, divacancies in 4H-SiC have garnered significant attention due to their long coherence time [2, 3], spin-selective optical transitions [4], and telecom-wavelength fluorescence [5]. Photoluminescence (PL) centers, specifically PL1-8 [6, 7], have been reported in 4H-SiC. In particular, PL5 and PL6 have similar optical and spin characteristics to those of PL1–4 [8, 6], while exhibiting superior charge stability under optical excitation [9, 10]. Experiment [8] and first-principles calculations [11] have unambiguously identified PL1–4 as the hh, kk, hk, and kh divacancies (), where the symbols refer to the hexagonal (h) or cubic (k) crystallographic sites of the vacancy pair. However, the PL5-8 centers have not yet been conclusively identified. A schematic diagram illustrating the 4H-SiC stacking and representative vacancy complexes is shown in Fig. 1(c).
PL5 and PL6 exhibit properties highly similar to the PL1–4 divacancy defects and are generally considered to be divacancies perturbed by their local environment. Initial hypotheses proposed that PL5 and PL6 correspond to divacancies (specifically kh and kk types) located near stacking faults [12]. However, this model was challenged by subsequent experiments [13], fueling ongoing debate and further investigation [14, 15]. Divacancies are typically generated via post-irradiation annealing [16] of SiC samples with a multifaceted point defect landscape, including intrinsic vacancies, antisite defects, and potential residual impurities. Given this complex defect landscape, an alternative plausible hypothes emerges: PL5 and PL6 may be complexes formed by combining a divacancy paired with another simple point defect nearby. Two prominent recent proposals advance this line of thought. One identifies PL5 as a kh divacancy complexed with a nearby carbon antisite () [17], referred to here as the DV-antisite model. The other, known as the OV model, attributes both PL5 and PL6 to oxygen-related defect complexes () [18]. Analogous to the divacancy notation, the OV configurations are differentiated by the hexagonal (h) or cubic (k) sites occupied by the and . Fig. 1(c) illustrates one such configuration: (hk). Under the widely held assumption that PL5 and PL6 share a common origin, a critical question arises: which scenario, DV-antisite or OV, provides a more robust explanation for PL6?
In this work, we systematically investigate the DV-antisite scenario for PL6 to identify its most probable structural origins. Using first-principles screening, we first analyze configurations based on zero-field splitting (ZFS), Kohn-Sham (KS) gap (used as an indicator of optical transition energy), and binding energy. This screening identifies two structures as promising candidates for PL6 within the DV-antisite framework. To further assess their validity, we compare these candidates with the leading OV(hh) structure proposed in Ref. [18], focusing on hyperfine interaction signatures. Our analysis suggests that the OV(hh) structure is the most probable origin of PL6. Additionally, we examine whether PL5 could be attributed to the proposed OV(hk) candidate, based on the hypothesis that PL5 and PL6 share a similar origin. However, inconsistencies between the calculated and measured ZFS parameter rule out the possibility that PL5 is a simple OV(hk) structure.
The remainder of this paper is organized as follows. Sec. II outlines our comprehensive first-principles computational methodology. This section details the general calculation setup (Sec. II.1), our strategy for identifying and screening potential defect candidates for PL6 under the DV-antisite hypothesis (Sec. II.2 and Sec. II.3), and the approach used to calculate the hyperfine interactions for the subsequent validation (Sec. II.4). Sec. III presents our findings. We first discuss the results of the screening within the DV-antisite framework to identify promising candidates in Sec. III.1. This is followed by a critical validation step, comparing the hyperfine interaction signatures of DV-antisite candidates and alternative OV model structures against experimental data to determine the most probable origin of PL6 (Sec. III.2). We then extend our discussion to consider implications for the PL5 center (Sec. III.3). Finally, Sec. IV summarizes our conclusions.
@@ II Methodology
Our investigation into the origin of PL6 commences with a systematic evaluation of the DV-antisite hypothesis, where PL6 is modeled as a divacancy paired with a nearby antisite. While the primary focus of the screening methodology detailed in this section is to identify promising candidates within this DV-antisite framework, the underlying first-principles calculation setup and characterization techniques are generally applicable and were consistently employed for all defect structures considered in this study, including the OV model. Thus, we begin by detailing the computational approaches for screening and characterizing these divacancy-antisite configurations.
@@ II.1 Calculation setup
II.2 Determining Candidate Complex Defects
To identify candidates for PL6 within the DV-antisite framework, we first consider experimental constraints. Experimental observations show that the rhombic ZFS parameter of PL6 is, in essence, zero within experimental uncertainties [6, 8], indicating that the principal axis of PL6 aligns with the c-axis. Consistent with this, angle-resolved measurements also indicate a -like symmetry for PL6 [13]. Furthermore, crucial constraints arise from PL6’s electronic and spin properties. PL6 exhibits an ground state. This state, similar to PL1–4, implies an even electron number and an electronic configuration resembling that of divacancies. Consequently, our search focuses on defect complexes that preserve these features. Introducing an additional vacancy is excluded as it would lead to . While neutral antisite defects act as isovalent substituents and maintain a comparable electronic environment when combined with a diavacancy, thereby preserving the fundamental spin and optical properties, highly charged states like are excluded due to their significant deviation from the divacancy’s electronic structure. Therefore, this study exclusively considers neutral hh/kk-type defect complexes incorporating neutral antisite defects.
With these constraints in mind, we proceed to search for possible complex defect structures. To explore variants of hh and kk configurations, we construct a cubic-like supercell with basis vectors , where are the primitive translation vectors of the primitive cell. In the first stage, a hh or kk defect was generated. Then, a cylinder was defined centered at this defect, with a height of 8 Å and a base radius of 7 Å, as the region to search for potential antisite locations. To reduce reduncancy, we leveraged the symmetry of the parent hh and kk divacancies to exclude symmetry-equivalent antisite locations within the cylinder. The final selection space consists of 116 symmetry-inequivalent antisite placements relative to the neutral divacancy: 56 for hh and 60 for kk. This broad configuration space is designed is capture all antisite placements with meaningful impact on the divacancy properties, considering the anisotropic spatial decay of spin density from the hh/kk center (see Fig. LABEL:spin-density in the Supplemental Material [26]).
@@ II.3 Screening Methodology for the PL6 Center Identification
To systematically identify the most plausible microscopic origin of PL6 among the candidate divacancy-antisite complexes, we established a multi-criteria screening framework integrating spin, optical, and thermodynamic properties. The selection criteria were designed to balance computational feasibility with experimental relevance, emphasizing three aspects: (1) consistency with experimentally measured ZFS parameters; (2) compatibility with photoluminescence characteristics through KS level analysis; (3) thermodynamic stability of defect complexes. This multi-property approach enables effective discrimination among thousands of possible defect configurations while maintaining manageable computational costs.
@@ II.3.1 ZFS Parameter Calibration and Error-Tolerant Thresholds
Note that experimentally measured values correspond to their absolute magnitudes. The ZFS parameters (, ) serve as critical fingerprints for spin defects. To address systematic errors in our DFT implementation, including PAW correction and spin decontamination errors, we performed benchmark calculations on the four well-characterized divacancies (PL1–4, i.e., hh, kk, hk, kh defects), as shown in Table 1. The calculated values are systematically overestimated by about 10%. For the values, our calculations exhibit a discrepancy of roughly when compared to the experimental data. Notably, the value for a strictly hh or kk divacancy should be zero; however, a small nonzero value emerges due to the absence of symmetry constraints in spin-density calculations. To ensure comprehensive coverage of potential PL6 candidates while accounting for these errors, we adopted the following conservative thresholds:
[1.365, 1.706] GHz ( times experimental = 1.365 GHz)
20 MHz (allowing a MHz deviation from over the experimental value of MHz)
@@
| Divacancies | (GHz) | (MHz) | ||
| expt | calc | expt | calc | |
@@ II.3.2 KS Level Difference as ZPL Predictor in Defect Screening
The photoluminescence characteristics provide critical complementary signatures for defect identification. For the symmetry divacancies, the ZPL originates from the spin-allowed transition [28], corresponding to two specific occupations of KS orbitals differed by promoting an electron from the highest occupied orbital to the lowest unoccupied orbital in the minority spin channel. The calculation of ZPL of PL1–4 has been detailed and tested among supercell size, k-point mesh, and functional selection [11]. However, such calculations are computationally demanding (e.g., HSE06 functional, a -only calculation with a supercell of 2400 atoms, or a calculation with -mesh for a 96-atom supercell) [11, 29].
The observed 0.1 eV blueshift in ZPL of PL6 over the hh/kk divacancies [8, 9] suggests a perturbation of the splitting between KS orbitals from an antisite defect. We establish a computationally efficient proxy for ZPL shifts through KS level analysis: the energy difference between the highest occupied () and lowest unoccupied () minority spin states systematically correlates with experimental ZPL trends in benchmarked defects (PL1–4). This relationship arises because ZPL energies and KS splitting respond proportionally to the perturbation. By requiring candidate complexes to exhibit enhanced KS splitting compared to isolated hh/kk centers, we effectively screen for structures capable of reproducing PL6’s characteristic ZPL shift.
@@ II.3.3 Binding Energy for Stability of Complex Defects
where is the formation energy of point defects. Negative values indicate exothermic complex formation, serving as a necessary (though not sufficient) condition for stability. While full thermodynamical analysis would require considering formation pathways [31] and configurational entropy [32], our screening conservatively requires to filter energetically favorable complexes.
At this stage, by applying criteria derived from several key experimental observations for the PL6 center—namely, its ZFS parameters ( GHz and MHz) and its ZPL blueshift relative to hh/kk divacancies—along with a theoretical requirement for thermodynamic stability, we aim to identify promising candidates. Specifically, the applied criteria are: D within GHz, within MHz, KS gap exceeding the reference, and negative binding energy. These are applied to the generated divacancy-antisite configurations under the DV-antisite hypothesis. Despite the intentionally permissive nature of these initial criteria, the geometric and electronic constraints inherent to these specific complex defects yield only a few viable configurations, demonstrating remarkable inherent selectivity in defect screening.
II.4 Hyperfine Interaction Calculation
To further validate the candidates emerging from the screening process and to critically compare them against alternative models, we calculate their hyperfine interaction signatures. The hyperfine interation is described by a tensor comprising isotropic part (Fermi contact, ) and anisotropic () components:
To interpret these calculations in the context of experimental data and the specific nature of our candidate defects, several considerations are pertinent. For the DV-antisite candidates involving a antisite, the substitution of a Si atom with a C atom means that if this antisite nucleus significantly contributes to the hyperfine spectrum, the opposite sign of the gyromagnetic ratio (compared to ) would lead to an inversion of the hyperfine coupling tensor’s sign at that particular site. In the case of the OV models, while the isotope possesses a nuclear spin, its extremely low natural abundance indicates that any associated hyperfine signals are too weak for detection in standard experiments without isotopic enrichment, making and the primary contributors to observable hyperfine structure. While the sign of the hyperfine interaction can, in principle, be determined by more complex methods like electron-nuclear double resonance (ENDOR) [34], the crucial experimental report that identified the hyperfine signatures of PL6 utilized ODMR, which resolves the magnitude of the splittings rather than the sign [35]. Therefore, to perform the most direct and rigorous comparison with the available experimental findings, our analysis focuses on the principal component , where is the unit vector along the c-axis, coinciding with the principal axis of the ZFS tensor for the -like centers.
@@ III Results and Discussion
III.1 Screening Potential PL6 Structures in DV-antisite Configurations
Applying the multi-criteria screening protocol (detailed in Sec. II.2) to the set of generated divacancy-antisite defect configurations, we effectively narrowed the vast configuration space. This process isolated only 12 structures consistent with the primary experimental constraints for PL6, as listed in Table 2. To systematically characterize these surviving configurations, we label the unperturbed divacancy and associated antisite defects using a grid-based positional encoding scheme. The numerical indices following an antisite symbol specify its position relative to the of the divacancy using grid coordinates along the -, -, and -axes, where negative displacements are denoted by overlined numbers. These coordinate axes are defined according to a gridded supercell with grid step sizes of 1.540 Å, 0.889 Å, and 0.628 Å along each respective axis. Complete implementation details, along with visual examples, are provided in Supplemental Material [26].
@@ III.1 Screening Potential PL6 Structures in DV-antisite Configurations
A notable feature of the results is that most binding energies are small (on the order of eV or less). An exception is , which has a more significant binding energy of -0.204 eV. This reveals that for most candidates, the attraction between the divacancy and antisite defect is weak, leading to defect complexes with limited stability. Another feature is that most of the candidates possess an whose absolute value is in the range of 10 to 20 MHz. Considering the precise experimental value of MHz for PL6, we further refine the selection by applying a more stringent of MHz to identify candidates best matching this key experimental signature. This yields the following candidates: , , and . Additionally, the remaining candidates show weak KS gap enhancement relative to the original kk divacancy defect. Among these candidates, and exhibit the most enhanced KS level splitting. Considering all these aspects, and emerge as the most compelling candidates for the PL6 center within the configuration space of DV-antisite complexes.
@@
Apart from the difference in local environment between and , the dominant difference between them is the distance between the kk center and the antisite, which is 4.10 Å for and 6.00 Å for . The longer distance in leads to a weaker interaction between the kk and the antisite, which is consistent with its smaller binding energy. For these two defects, we further calculated the actual ZPL energy by using constrained DFT calculation and the 0/1- charge transition level by the PBEsol functional. While this set of computational parameters cannot reliably reproduce experimental ZPL values, it should still capture the relative shifts between different configurations. This allows for a comparative analysis of trends despite the absolute discrepancy. A comparison of the results with kk is shown in Table 3. Notably, exhibits a significant ZPL shift of 0.13 eV relative to kk, whereas the ZPL shift for is considerably smaller, only 0.03 eV. Moreover, the 0/-1 charge transition level of is closer to the conduction band minimum (CBM) than kk, consistent with the experimental charge dynamics model [9]. Thus, although has a relatively small binding energy, its ZPL shift and charge transition level characteristics establish it as a notable candidate for PL6 under the DV-antisite hypothesis. Although the is closer to the kk center in , the properties of the divacancy mainly arise from the dangling bond of the C atom, which originates from the inside the divacancy. Although anti-site in is farther from the kk center, it is closer to the dangling bond of the kk center, which is the reason why the impact of the antisite defect on the kk center is larger in than in .
To summarize,, we utilized experimental data to screen potential PL6 candidates, identifying the most likely structures under the DV-antisite scenario. However, there are also many other defects considered in our calculation that serve as good spin defects but are not PL6 candidates. They are listed in Table LABEL:tab:other_spin_defects. Some discussion on them is also given in the Supplemental Material [26].